OpenStudy (anonymous):

A cyclist accelerates from 0m/s to 8m/s in 3 seconds. What is his acceleration? Is this acceleration higher than that of a car which accelerates from 0 to 30 m/s in 8 seconds?

OpenStudy (queelius):

Assuming constant acceleration, then v(t) = a*t

OpenStudy (queelius):

We are given that v(3) = 8.

OpenStudy (queelius):

Thus, a*3 = 8. Solve for a?

OpenStudy (anonymous):

Assuming acceleration is constant, acceleration of cyclist = ( V(f) - V(i) ) / t = (8-0)/3 = 2.67 ms-2 acceleration of car = (30-0)/8 = 3.75 ms-2 Hence, acceleration of cyclist is not higher than that of car.